QUESTION 18 A 2kg steel ball hangs on a 8.0-m-long string. The ball is pulled sideways to an angle of 45°, then released. At(6 5 points) What is the compression in the spring (m)?
Answer #1

a It mass mitme ОХ 8(1-ros of 0-094 = cosa=0-9883 It is elastic collision of the second is initially at at rust, so- V = mi-m